What are the dimesions of length in


force*discplacemrnt/time


Options-


1. 2


2. -2


3. 0


4. None of these


Please explain in detail

Grade:10

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Amit

force= mass X acc

acc= L/T2

force= ML/T2

Displacement=L

given expression will be

(ML/T2)*L/T

ML2T-3

Hence answer is 2 that is first option

 

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

vikas askiitian expert
13 years ago

C = force*displacement/time

F=MLT-2

DISPLACEMENT = L

TIME = T

dimensional formula of C is MLT-2 L/T =ML2T-3

DIMENSION OF L IS 2

ankur kalra
13 years ago

ans is 2